

********************************************************************
UARS WEEKLY STATUS REPORT - 16 JUN 2000
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM continues to gather a near normal set of daily solar scans
scheduled for periods of predicted telemetry.
Despite the unexpected spacecraft power off on June 7, the monthly
lamp calibration was completed successfully beforehand.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM spectrometer daylight operation was interrupted beginning June 5
due to the power negative condition which existed with the UARS batteries.
PEM Daylight operations resumed on June 14. The delay from activation on
June 9 was due to battery low conditions still occurring. The PEM spectrometer
operation in daylight is active under PMON control* when PMON is enabled.
At the present time, PEM energy deposition is not being generated.
Recent data configuration changes have affected PEM software causing multiple
failures. These failures are being reviewed by the PEM team. In order to
correct for new telemetry requirements, PEM will need to update processing
code with a patch delivery. The delivery patch will not include changes which
would allow the PEM software to execute on a DEC Alpha machine.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII is scheduled to operate June 18 - 26. It will be configured
for the observation of polar mesospheric clouds.
***********
MPG
***********
From an action item of the UARS Mission Status Review meeting (on
June 8) the MPG has queried instrument teams about their use of
UARS Quick Look data. Responses have been received from three
instruments so far, and these three do not use Quick Look data.
Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to assist the Project Scientist in
scheduling of instrument operations.
***********
CDHF
***********
STATUS:
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
- All requested reprocessing is complete as of this reporting period.
CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, HRDI,
MLS, SUSIM, Solstice, and WINDII.
- Waiting validation results from the HRDI group on successful
production testing done on the Rehost System.
- System tests being done.
***********
FOT
***********
REPORTING PERIOD
This report covers 03 June 2000 (Orbit 47703, GMT Day 155) through 09 June 2000
(Orbit 47807, GMT Day 161).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 157/1706 47743 Off
HRDI 157/2138 47745 Standby
SOLSTICE 159/2000 47774 On
ACRIM 160/0105 47778 On
SUSIM 160/1346 47786 On
HRDI 161/0007 47792 Restart
HRDI 161/0331 47794 Day Only
HALOE 161/1306 47800 On
A Forward-to-Reverse Yaw Around maneuver (number 89) was successfully executed
in Orbit 47742 beginning 157/16:01:29 and completing at 157/16:38:39 on 05 June
2000. The SSPP Instruments (ACRIM, SOLSTICE, SUSIM) were enabled as well as
HRDI and PEM in day only operations. Due to the SA position at 269 degrees,
the effective night length of the spacecraft after the maneuver was
approximately 55 minutes. Post maneuver analysis of data indicate that during
the Yaw Around, the end-of-night load bus voltage was 23.2 volts and the
average depths of discharge (DODs) on Batteries 2 and 3 were approximately 24.5
and 24.6 percent respectively. The next (reverse-to-forward) YAW is scheduled
for 04 July 2000 (DOY 186).
The beta angle ranged between the angles of 11.7 degrees to 13.7 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
34.0 degrees on 20 June 2000 (DOY 172).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.2 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.6 to 1.0 degrees C. Additional battery performance data for each
battery on 02 June 2000 and 09 June 2000 are provided in Appendix D.
The clock error ranged from -13.8 to 12.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 13
% ALERTS = 02
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 07
% SUSIM = 04
% HRDI = 04
% HALOE = 02
% PEM = 01
% SOLSTICE = 01
% ACRIM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
Five (5) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/145 - 2287.497008
- Transponder B: 00/152 - 2287.498920
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/160 show the observatory data loss to be 296 hours, 20
minutes, 7.713 seconds (an increase of 40 minutes 13.672 seconds since the last
report period). This is a 0.40 percent data loss which equals a 99.60 percent
data capture for the mission.
The increase of 40 minutes 13.672 seconds occurred over two (2) days:
- 0 minutes 08.576 seconds on Day 156 due to a data hit.
- 40 minutes 05.096 seconds on Day 159 due to a late AOS and an early off to
conserve power.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The OBC Telemetry Monitors (TMONs) 1 and 2 tripped during this report period.
These TMONs shed load when the EON Voltage on Batteries one and two get too
low. An investigation has begun to find the cause of this anomaly.
Preliminary findings suggest an insufficient power margin caused the TMONs to
trip, not a battery degradation problem. All instruments have been returned to
their scheduled operations with the exception of PEM, which should resume its
operations soon.
The FOT continues to test the new TMONs on the simulator. Once the TMONs have
been thoroughly tested, they will be presented for review and approval before
uplink.
The FOT has begun writing the online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure will
be tested on the simulator and put online for immediate execution should an
anomalous condition take place.
MEETINGS
The FOT attended the UARS Monthly Meeting on 07 June 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for July 04, 2000 (DOY 186). This maneuver
will be a Reverse to Forward Flight maneuver (Southern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 09 June 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
47708 155/1000 00-249 OBC Clk Adj fr 20.6 to 0.00 msec/day
47713 155/1832 00-250 SUSIM SUNACTIV
47714 155/1939 00-251 OBC Clk Adj fr 0.00 to -20.6 msec/day
47718 156/0223 00-252 SUSIM Command request
47723 156/1002 00-253 SUSIM Cmds to replace daily microload
47734 157/0305 00-254 OBC Clk Adj fr -20.6 to 0.00 msec/day
47741 157/1506 00-255 OBC YAW AROUND
47743 157/1706 00-256 PEM Power OFF DUE TO S/C LOW PWR
47745 157/2138 00-257 HRDI HRDI TO STBY
47773 159/1819 00-257 HRDI Power ON to STBY only
47774 159/2000 00-258 SOLSTICE Power ON
47775 159/2149 00-259 NBTR PLAYBACK
47778 160/0105 00-260 ACRIM Power ON
47786 160/1346 00-261 SUSIM Power ON
47792 161/0007 00-257 HRDI RESTART
47794 161/0331 00-262 HRDI HRDI: DAY-OPS ONLY
47796 161/0625 00-263 NBTR NOOPS D-161
47800 161/1306 00-264 HALOE Power ON: P-1
47804 161/1904 00-265 NBTR NOOPS D-162
47805 161/2146 00-266 HALOE Power ON: P-2
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 09 June 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI PSE QB CURRENT MON 47724-30
HRPTSPBIMON R-LO PTS PB CURRENT MON 47724
HRPSEQBIMON Y-HI PSE QB CURRENT MON 47740-3
OBSVLBEON Y-LO EON LOAD BUS VOLTAGE 47741-4
OBSBATVEON2 Y-LO EON Bat 2 voltage 47741-4
OBSBATVEON3 Y-LO EON Bat 3 voltage 47741-4
PWLBV Y-LO LOAD BUS VOLTAGE 47742-3
HRPTSPBIMON R-HI HRDI PB CURRENT 47746
HRPTSPBIMON Y-HI HRDI PB CURRENT 47746
HRPSEQBIMON Y-HI PSE QB CURRENT 47754-9
PWBAT3V Y-LO BATT 3 VOLTAGE 47742-3
PWBAT2V Y-LO BATT 3 VOLTAGE 47742-3
CDUNRG28 Y-LO 28 VOLT UNREG BUS 47742-3
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- - ----------- ------
SOPUPERSTATA Solstice voltage low 47766
ACINVMCNREDL ACRIM volt is red low 47766
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 09 June 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
47708 155/10:02 (03 June 00) from 20.6 to 0.00 msec/day
47714 155/19:47 (03 June 00) from 0.00 to -20.6 msec/day
47734 157/03:08 (05 June 00) from -20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 09 June 2000
02 June 2000 - GMT Day 154 Beta = 11.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ----------- ------ ------- ------
1 +22.4/+0.0 -3.05/-4.62 +0.0/+0.0 15.7 0.000 0.0
2 +39.2/-22.4 +6.44/+3.87 +30.4/-14.4 24.5 1.040 20.4
3 +22.4/-67.2 +5.15/+2.92 +29.6/-14.4 24.3 1.038 19.7
09 June 2000 - GMT Day 161 Beta = 13.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------- ------- ------
1 +33.6/+11.2 -3.05/-4.62 +0.0/+0.0 15.7 0.000 0.0
2 +44.8/-22.4 +5.79/+4.51 +29.6/-16.0 24.3 1.030 20.6
3 +22.4/-67.2 +4.83/+3.23 +28.8/-15.6 24.2 1.027 20.2
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 09 June 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-024 FOT 47742 157/1635 N/A TMON 1 trip
00-025 FOT 47766 159/0657 N/A TMONS 1 & 2 tripped
00-026 FOT 47770 159/1300 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-007 FOT see list May 00 14170 Generic Late Acq.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-024 FOT 47742 157/1635 N/A TMON 1 trip
00-025 FOT 47766 159/0657 N/A TMONS 1 & 2 tripped
00-026 FOT see list June 00 14170 Generic Late Acq.
SEVNTFW00/23